# Large-scale distributed synchronization systems, using a cancel-on-completion redundancy mechanism

### v1

@@ 3.5 Steady-state asymptotic independence definition
(i) We say that steady-state asymptotic independence (SSAI) holds for speed $v>0$ for the anchored system, which is either left-regulated, right-regulated or both-side-regulated, if Markov process $x^{n}(\cdot)$ is stable (positive recurrent) for all large $n$ and
@@
(ii) We say that steady-state asymptotic independence (SSAI) holds for the unregulated (free) system, if centered Markov process $x^{n}(\cdot)$ is stable (positive recurrent) for all large $n$ and
@@
For the anchored system with frame $[0,\infty)$ (i.e. left-regulated at 00 and unregulated on the right), the following property will be called SSAI for the full range of loads (SSAI-FRL): there exists $\underline{v}$ such that for any $v>\underline{v}$ the convergence $x^{n}(\infty)\Rightarrow x^{*,v}$ holds for some fixed $x^{*,v}\in\tilde{\cal X}^{pr}$ and $x^{*,v}_{0}\uparrow 1$ as $v\downarrow\underline{v}$. The name SSAI-FRL is natural, because in terms of the multi-server queueing system $x^{*,v}_{0}$ is the limiting fraction of occupied servers – the system load.
@@ 4.2 Reduced systems and corresponding job subclasses
We now define reduced systems and job subclasses – the notions used in the analysis.
Consider a given system, with fixed $n$, with the set of job classes $\{j\}$ and corresponding parameters $\{k_{j},d_{j},F_{j},\sigma_{j}\}$ for each $j$. (We will not use superscript $n$ in this subsection to avoid clogging notation.) Let two parameters, $\epsilon\geq 0$ and $\delta\geq 0$, with $\epsilon+\delta<1$, be fixed. Suppose that fraction $\epsilon$ of all particles are located at +$\infty$+ ∞ (and therefore alway remain there) and fraction $\delta$ of all particles are located at $\infty$- ∞ (and therefore alway remain there). The remaining fraction $1-\epsilon-\delta$ of particles are “regular” – they are initially (and then always) at located at points in $\mathbb{R}$. What is the evolution of the “standard” particles? It can be equivalently described as the evolution of the following $(\epsilon,\delta)$-reduced system. First of all, in the $(\epsilon,\delta)$-reduced system there are $(1-\epsilon-\delta)n$ particles, instead of $n$. For each class $j$ in the original system, there are multiple classes in the reduced system. Specifically, consider two fixed integers $\ell$ and $m$, such that $0\leq\ell<k_{j}$ and $\ell<m\leq d_{j}$. If a class $j$ job arrives into the original system, and it happens that it selects exactly $\ell$$\infty$- ∞-particles” and exactly $d_{j}-m$+$\infty$+ ∞-particles,” then, in terms of the reduced system, this is equivalent to the arrival of a job of the class $j^{m,\ell}$ with the following parameters (we will use notation $\omega=j^{m,\ell}$ for brevity): $k_{\omega}=k_{j}-\ell$, $d_{\omega}=m-\ell$; the distribution $F_{\omega}$ of the component sizes is the projection of $F_{j}$ on the distribution of first $m$ components; the arrival rate (per server) is
$\sigma_{\omega}=\sigma_{j}\mathbb{P}\{\mbox{a job $j$ selects exactly $\ell$ ``$-\infty$-particles'' and exactly $d_{j}-m$ ``$+\infty$-particles''}\}/(1-\epsilon-\delta).$ (8)
Indeed, from the point of view of reduced system, the selected $d_{j}-m$+$\infty$+ ∞-particles” can be ignored, and the selected $\ell$$\infty$- ∞-particles” reduce the number of particles whose movement needs to be completed from $k_{j}$ to $k_{\omega}=k_{j}-\ell$; thus, in the reduced system we only need to consider the selection of $m-\ell$ particles, of which the movement of $k_{j}-\ell$ needs completion. The renormalization by factor $(1-\epsilon-\delta)^{-1}$ in (8) is needed because the number of particles is $(1-\epsilon-\delta)n$ instead of $n$. The constraint $\ell<k_{j}$ is because if $k_{j}$ or more “$\infty$- ∞-particles” are selected by a class $j$ in the original system, this has no impact on the reduced system; similarly, the constraint $\ell<m$ is because, otherwise, no particle in the reduced system is selected.
@@ 5 Mean-field limits
If $x$ is an ML-FP for the anchored left-regulated system with frame $[A,\infty)$, the value $x_{A}$ at the left boundary will be called the load of $x$ (because this value corresponds to the fraction of occupied server in the multi-server queueing system).
@@
(b) $x_{w}$ is non-increasing $C$-Lipschitz, with constant $C$ dependent only on the system parameters, in the interval $[A,\infty)$.
(c) For almost all $w\in(A,B)$ w.r.t. Lebesgue measure, , we have
$-v\frac{d}{dw}x_{w}=\lambda h(x_{(-\infty,w]}).$ (14)
@@
For a given anchored system, let us define the following operator $\cal A$ which maps $\tilde{\cal X}$ into itself. Let a distribution (“environment”) $x\in\tilde{\cal X}$ be fixed; $x$ may or may not be proper. Consider a single particle moving in the frame $[A,B]$ (where the left boundary $A$ may be at $\infty$- ∞, and right boundary $B$ may be at +$\infty$+ ∞) as follows. Unless and until the particle jumps, it moves left at speed $v$; if/when the particle “hits” the left boundary $A$, it stays there until next jump; there is a rate $\alpha$ Poisson process of time points when the the particle jumps forward; if the jump occurs when the particle location is $w$, the jump size distribution is $J^{(x,w)}(\cdot)$, independent of the process history (besides current location $w$); however, if the particle would jump over point $B$, it lands exactly at $B$ instead. Then ${\cal A}x\in\tilde{\cal X}$ is the stationary distribution of the particle. Note that, if a particle is located at $\infty$- ∞ or +$\infty$+ ∞, it stays there forever. Consequently, $({\cal A}x)_{\infty}=x_{\infty}$ and $({\cal A}x)_{-\infty}=x_{-\infty}$.
@@
Consider ML-FPs in anchored systems. Suppose we have a sequence, indexed by $\ell\to\infty$, of an anchored system parameters $v^{(\ell)}>0$, $\{\sigma_{j}^{(\ell)}\}$, $\{F_{j}^{(\ell)}\}$, $A^{(\ell)}$, $B^{(\ell)}$, converging, as $\ell\to\infty$, to the limiting set of parameters $v>0$, $\{\sigma_{j}\}$, $\{F_{j}\}$, $A$, $B$, with $-\infty\leq A<B\leq\infty$. (Index $\ell$ is not the number $n$ of particles in a pre-limit system. Instead, it indexes ML-FPs for different parameter sets. Also, $A^{(\ell)}$, $B^{(\ell)}$ may be finite or infinite.) Suppose, for each $\ell$ there exists an ML-FP $x^{(\ell)}\in\tilde{\cal X}$ (proper or not proper on either side) for the corresponding parameter set, and $x^{(\ell)}\to x\in\tilde{\cal X}$. Then $x$ is an ML-FP for the limiting set of parameters.
@@
Consider any ML-FP (=DE-FP) $x=(x_{w})\neq x^{**}$, for an anchored system, with left and right boundaries (if any) at $-\infty\leq A<B\leq\infty$. (Note that, if $A$ is finite, then necessarily, $x_{w}<1$.) Then, $x$ satisfies the following additional properties. Function $x_{w}$ is strictly decreasing in $[A,B)$. The right derivative $x^{\prime+}_{w}$ exists at every point $w\in[A,B)$. The left derivative $x^{\prime-}_{w}$ exists, and $x^{\prime-}_{w}\leq x^{\prime+}_{w}$, at every point $w\in(A,B)$. (These derivatives are non-positive.) Equation (14) holds for the right derivative $x^{\prime+}_{w}$ at every point $w\in[A,B)$. The right derivative $x^{\prime+}_{w}$ is RCLL, the left derivative $x^{\prime-}_{w}$ is LCRL, and each is negative, bounded away from 00 on any finite interval $[A\wedge a,b]$, where $-\infty<a<b\leq B$.
@@
(i) Consider the anchored left-regulated system, with frame $[0,\infty)$. Consider the ML $x(\cdot)$, starting with the “empty” initial state $x(0)=x^{\emptyset}$. Then $x(t)$ is monotone non-decreasing and $x(t)\to x^{*}$, where $x^{*}$ is an ML-FP, and moreover it is the minimum ML-FP. (That is, $x^{*}\leq x$ for any ML-FP $x$.) If $x^{*}\neq x^{**}$, then $x^{*}$ is proper.
(ii) Consider the anchored right-regulated system, with frame $(-\infty,0]$. Consider the ML $x(\cdot)$, starting with the “empty” initial state $x(0)=x^{\emptyset}$. Then $x(t)$ is monotone non-increasing and $x(t)\to x^{*}$, where $x^{*}$ is an ML-FP, and moreover it is the maximum ML-FP. (That is, $x^{*}\geq x$ for any ML-FP $x$.) If $x^{*}\neq 1-x^{**}$, then $x^{*}$ is proper. (ML state $1-x^{**}$ corresponds to the entire particle mass at $\infty$- ∞.)
@@
For any anchored two-side-regulated system, there exists a unique ML-FP $x$ for any speed $v>0$, and the dependence on $v$ is continuous.
@@ 6 Preliminary definitions and results (generalizing those in [15]), related to SSAI
For a given $n$ consider Markov process $x^{n}(\cdot)$ for an anchored system, and the corresponding centered process $\mathring{x}^{n}(\cdot)$. Process $\mathring{x}^{n}(\cdot)$ is not Markov, because the information about the location of the boundaries is “lost” after centering; it is a projection of Markov process $\mathring{x}^{n}(\cdot)$.
@@
Consider the anchored left-regulated system with frame $[0,\infty)$. Suppose Assumptions 3.2, 3.3 and 3.4 hold. Then there exist $\bar{C}>0$ and $\bar{n}$ such that, uniformly in $n\geq\bar{n}$ and corresponding (for each $n$) speeds $v$ such that the process $x^{n}(\cdot)$ is stable, we have
@@
Suppose there exists a class $j$, $0\leq m_{j}<k_{j}$, and $\epsilon>0$, such that the component size distribution $F_{j}$ dominates a scaled (by a positive constant) measure such that $m_{j}$ components are zero and the remaining $d_{j}-m_{j}$ components distribution is the Lebesgue measure on $[0,\epsilon]^{d_{j}-m_{j}}$. Then Assumption 6.2 holds.
@@
If process $\mathring{x}^{n}(\cdot)$ in a free system is positive recurrent, then the average velocity $v_{n}$ of the unanchored free particle process $x^{n}(\cdot)$ is well-defined, with multiple possible equivalent definitions. For example,
[…]
@@
Consider an anchored left-regulated system, with frame $[0,\infty)$. We say that, in the anchored left-regulated system with frame $[0,\infty)$, this condition holds for speed $v>0$, if for any sufficiently small $b\geq 0$, there is at most one DE-FP $x$ with $x_{\infty}=b$.
@@
Consider the left-regulated system with the frame $[0,\infty)$. Suppose Assumptions 3.2, 3.3 and 3.4 hold. Consider our asymptotic regime, with the number of particles $n\to\infty$, such that for all large $n$ the process is stable and, moreover, $\rho^{n}=\mathbb{E}x^{n}_{0}(\infty)\to\rho<1$. Suppose that Condition 6.5 holds for speed $v>0$. Then SSAI holds for this $v$, i.e. $x^{n}(\infty)\Rightarrow x^{*}$, where $x^{*}$ is the ML-MFP and it is such that $x^{*}_{0}=\rho$.
@@
The proof of Lemma 6.8 is straightforward. Consider any finite locations of $m-\ell$ particles in the reduced system; consider original system with $m-\ell$ particles located as in the reduced system, $\ell$ particles located at $-C$, $d_{j}-m$ particles located at $C$; let $C\uparrow\infty$.
@@ 7 Main results
Consider ML-FPs for a family of anchored systems, with the following set of parameters taking values within a compact set: speed $v$; limiting arrival rates $\{\sigma_{j}\}$; and the distributions $\{F_{j}\}$ of the job component sizes (convergence of distributions is in the weak sense). Assume further that the family of parameters is such that: speed $v$ is bounded away from 00; Assumption 3.2 holds uniformly; Assumption 3.3 holds uniformly in that it holds for the same class $j$, with the corresponding probability bounded away from 00, and $\sigma_{j}$ bounded away from 00. Then, for some $\bar{C}>0$, uniformly in that family of parameters and uniformly across all free and regulated systems, an ML-FP $x$ is such that $\Phi_{1}(\mathring{x})\leq\bar{C}<\infty.$
@@
Proof. For a two-side-regulated system, with frame $[0,B]$ there exists a unique ML-FP for any speed $v>0$. For a sequence of frame sizes $B\to\infty$, choose drift $v(B)$, such that the median of the fixed point is exactly in the middle, at point $B/2$. Let $x^{(B)}$ be such ML-FP shifted left by $B/2$, so that the median is at 00. Let us choose a subsequence of $B$ along which $v(B)\to v$. Limiting speed $v$ cannot be 00 or +$\infty$+ ∞, because otherwise the median could not stay in the middle. (Indeed, if the limiting speed is infinity, then the probability mass of the left boundary has to stay bounded away from 00 – otherwise the average drift to the right cannot match the average drift to the left; in fact, that mass goes to $1$, so the distribution concentrates on the left boundary. If the limiting speed is zero, then the distribution, in a weak limit, must concentrate at the right boundary – otherwise the average drift to the right cannot go to 00 to match the average drift to the left.) Then, by Theorem 7.1, $\Phi_{1}(\mathring{x}^{(B)})\leq C<\infty$, uniformly in all $c$. This in turn implies that $\Phi_{1}(x^{(B)})\leq C<\infty$, uniformly in all $B$, because $|\bar{x}^{(B)}|$ must stay bounded. So, we can choose a further subsequence of $B$ along which $x^{(B)}\to x$, where $x$ is proper. Then, by Lemma 5.12, $x$ is a proper free ML-FP with speed $v$. $\Box$
@@
[…]
@@
[…]
Note that for any $10\leq\epsilon<$ 1,
@@
and $\chi^{(0)}$ is nothing else, but the stationary distribution of a particle moving in $(-\infty,0]$, with the rate $\alpha$ Poisson process of independent jumps to the right (but not over the right boundary at 00), with size distribution $H(\cdot)$, and moving left at speed $v$ between the jumps. We conclude that $\chi^{(\epsilon)}$ is the unique ML-FP (=ML-MFP) of the artificial $(\epsilon,0)$-reduced system with the right boundary at 00.
As will be seen in the proof of Theorem 7.3, $\chi^{(\epsilon)}$ gives the asymptotic shape of the left tail of any DE-FP with speed $v$. Informally, this is because, when $1-\epsilon$ is close to 00, the $(\epsilon,0)$-reduced original system is approximately the $(\epsilon,0)$-reduced artificial system.
@@
(i) In the anchored system, for any $v$ satisfying (24), there exists the unique DE-FP with the proper left tail. This solution will either (a) hit the $w$-axis or (b) will converge to some $\epsilon_{*}\geq 0$ as $w\uparrow\infty$. In case (a): for any $10\leq\epsilon<$ 1 there is a unique ML-FP with proper left tail for the $(\epsilon,0)$-reduced right-regulated system. In case (b): for any $\epsilon_{*}<\epsilon<1$ there is a unique ML-FP with proper left tail for the $(\epsilon,0)$-reduced right-regulated system; for $\epsilon=\epsilon_{*}$ there is a unique ML-FP with proper left and right tails for the $(\epsilon,0)$-reduced free system; for $\epsilon<\epsilon_{*}$ there is a unique ML-FP with proper left tail, but not proper on the right, for the $(\epsilon,0)$-reduced free system.
(ii) In the anchored system, for any $v$ satisfying (24), and any $10\leq\epsilon<$ 1, there exists at most one ML-FP with proper left tail for the $(\epsilon,0)$-reduced system with speed $v$; this ML-FP is either for the right-regulated system or for the free system, with proper or not proper right tail.
@@
(i) For the anchored free system, there exists a range $[v_{min},v_{max}]$ of speeds $v$, with $0<v_{min}\leq v_{max}<\infty$ for which a proper ML-FP $x$ exists.

### v2

@@ 3.5 Steady-state asymptotic independence definition
(i) We say that steady-state asymptotic independence (SSAI) holds for speed $v>0$ for the anchored system, which is either left-regulated, right-regulated or both-side-regulated, if the Markov process $x^{n}(\cdot)$ is stable (positive recurrent) for all large $n$ and
@@
(ii) We say that steady-state asymptotic independence (SSAI) holds for the unregulated (free) system, if the centered Markov process $\mathring{x}^{n}(\cdot)$ is stable (positive recurrent) for all large $n$ and
@@
For the anchored system with frame $0,\infty)$ (i.e. left-regulated at 0 and unregulated on the right), the following property will be called SSAI for the full range of loads (SSAI-FRL): there exists $\underline{v}$ such that for any $v>\underline{v}$ the convergence $x^{n}(\infty)\Rightarrow x^{*,v}$ holds for some fixed $x^{*,v}\in\tilde{\cal X}^{pr}$ and $x^{*,v}_{0}\uparrow 1$ as $v\downarrow\underline{v}$. The name SSAI-FRL is natural, because in terms of the multi-server queueing system $x^{*,v}_{0}$ is the limiting fraction of occupied servers – the system load. The notion of SSAI-FRL was introduced in [15]; see [15, section 1.2] for a discussion of this property is not automatic – and may be challenging to prove – for non-work-concerving systems.
@@ 4.2 Reduced systems and corresponding job subclasses
We now define reduced systems and job subclasses – the notions used in the analysis throughout the paper. One of the primary uses of the reduced system notion is as follows: it allows us to interpret a “truncation” of an ML fixed point (defined in Section 5) as an ML fixed point of a reduced system, which in turn allows us to use properties of the latter to analyze the original ML fixed point. (This is done, for example, in Theorem 7.3, as well as in other places.)
Consider a given system, with fixed $n$, with the set of job classes $\{j\}$ and corresponding parameters $\{k_{j},d_{j},F_{j},\sigma_{j}\}$ for each $j$. (We will not use superscript $n$ in this subsection to avoid clogging notation.) Let two parameters, $\epsilon\geq 0$ and $\delta\geq 0$, with $\epsilon+\delta<1$, be fixed. Suppose that fraction $\epsilon$ of all particles are located at $\infty$ (and therefore alway remain there) and fraction $\delta$ of all particles are located at -$\infty$ (and therefore alway remain there). The remaining fraction $1-\epsilon-\delta$ of particles are “regular” – they are initially (and then always) at located at points in $\mathbb{R}$. What is the evolution of the “regular” particles? It can be equivalently described as the evolution of the following $(\epsilon,\delta)$-reduced system. First of all, in the $(\epsilon,\delta)$-reduced system there are $(1-\epsilon-\delta)n$ particles, instead of $n$. For each class $j$ in the original system, there are multiple classes in the reduced system. Specifically, consider two fixed integers $\ell$ and $m$, such that $0\leq\ell<k_{j}$ and $\ell<m\leq d_{j}$. If a class $j$ job arrives into the original system, and it happens that it selects exactly $\ell$-$\infty$-particles” and exactly $d_{j}-m$$+\infty$-particles,” then, in terms of the reduced system, this is equivalent to the arrival of a job of the class $j^{m,\ell}$ with the following parameters (we will use notation $\omega=j^{m,\ell}$ for brevity): $k_{\omega}=k_{j}$-$\ell$, $d_{\omega}=m$-$\ell$; the distribution $F_{\omega}$ of the component sizes is the projection of $F_{j}$ on the distribution of first $m$-$\ell$ components; the arrival rate (per server) is
[…]
Indeed, from the point of view of reduced system, the selected $d_{j}-m$$+\infty$-particles” can be ignored, and the selected $\ell$-$\infty$-particles” reduce the number of particles whose movement needs to be completed from $k_{j}$ to $k_{\omega}=k_{j}$-$\ell$; thus, in the reduced system we only need to consider the selection of $m$-$\ell$ particles, of which the movement of $k_{j}$-$\ell$ needs completion. The renormalization by factor $(1-\epsilon-\delta)^{-1}$ in (8) is needed because the number of particles is $(1-\epsilon-\delta)n$ instead of $n$. The constraint $\ell<k_{j}$ is because if $k_{j}$ or more “-$\infty$-particles” are selected by a class $j$ in the original system, this has no impact on the reduced system; similarly, the constraint $\ell<m$ is because, otherwise, no particle in the reduced system is selected.
@@ 5 Mean-field limits
If $x$ is an ML-FP for the anchored left-regulated system with frame $A,\infty)$, the value $x_{A}$ at the left boundary will be called the load of $x$ (because this value corresponds to the fraction of occupied server in the multi-server queueing system).
@@
(b) $x_{w}$ is non-increasing $C$-Lipschitz, with constant $C$ dependent only on the system parameters, in the interval $A,\infty)$.
(c) For almost all $w\in(A,B)$ w.r.t. Lebesgue measure, we have
$v\frac{d}{dw}x_{w}=\lambda h(x_{(-\infty,w]}).$ (14)
@@
For a given anchored system, let us define the following operator $\cal A$ which maps $\tilde{\cal X}$ into itself. Let a distribution (“environment”) $x\in\tilde{\cal X}$ be fixed; $x$ may or may not be proper. Consider a single particle moving in the frame $[A,B]$ (where the left boundary $A$ may be at -$\infty$, and right boundary $B$ may be at $\infty$) as follows. Unless and until the particle jumps, it moves left at speed $v$; if/when the particle “hits” the left boundary $A$, it stays there until next jump; there is a rate $\alpha$ Poisson process of time points when the the particle jumps forward; if the jump occurs when the particle location is $w$, the jump size distribution is $J^{(x,w)}(\cdot)$, independent of the process history (besides current location $w$); however, if the particle would jump over point $B$, it lands exactly at $B$ instead. Then ${\cal A}x\in\tilde{\cal X}$ is the stationary distribution of the particle. Note that, if a particle is located at -$\infty$ or $\infty$, it stays there forever. Consequently, $({\cal A}x)_{\infty}=x_{\infty}$ and $({\cal A}x)_{-\infty}=x_{-\infty}$.
@@
Consider ML-FPs in anchored systems. Suppose we have a sequence, indexed by $\ell\to\infty$, of an anchored system parameters $v^{(\ell)}>0$, $\{\sigma_{j}^{(\ell)}\}$, $\{F_{j}^{(\ell)}\}$, $A^{(\ell)}$, $B^{(\ell)}$, converging, as $\ell\to\infty$, to the limiting set of parameters $v>0$, $\{\sigma_{j}\}$, $\{F_{j}\}$, $A$, $B$, with $\infty\leq A<B\leq\infty$. (Index $\ell$ is not the number $n$ of particles in a pre-limit system. Instead, it indexes ML-FPs for different parameter sets. Also, $A^{(\ell)}$, $B^{(\ell)}$ may be finite or infinite.) Suppose, for each $\ell$ there exists an ML-FP $x^{(\ell)}\in\tilde{\cal X}$ (proper or not proper on either side) for the corresponding parameter set, and $x^{(\ell)}\to x\in\tilde{\cal X}$. Then $x$ is an ML-FP for the limiting set of parameters.
@@
Consider any ML-FP (=DE-FP) $x=(x_{w})\neq x^{**}$, for an anchored system, with left and right boundaries (if any) at $\infty\leq A<B\leq\infty$. (Note that, if $A$ is finite, then necessarily, $x_{w}<1$.) Then, $x$ satisfies the following additional properties. Function $x_{w}$ is strictly decreasing in $[A,B)$. The right derivative $x^{\prime+}_{w}$ exists at every point $w\in[A,B)$. The left derivative $x^{\prime-}_{w}$ exists, and $x^{\prime-}_{w}\leq x^{\prime+}_{w}$, at every point $w\in(A,B)$. (These derivatives are non-positive.) Equation (14) holds for the right derivative $x^{\prime+}_{w}$ at every point $w\in[A,B)$. The right derivative $x^{\prime+}_{w}$ is RCLL, the left derivative $x^{\prime-}_{w}$ is LCRL, and each is negative, bounded away from 0 on any finite interval $A\wedge a,b]$, where $\infty<a<b\leq B$.
@@
(i) Consider the anchored left-regulated system, with frame $0,\infty)$. Consider the ML $x(\cdot)$, starting with the “empty” initial state $x(0)=x^{\emptyset}$. Then $x(t)$ is monotone non-decreasing and $x(t)\to x^{*}$, where $x^{*}$ is an ML-FP, and moreover it is the minimum ML-FP. (That is, $x^{*}\leq x$ for any ML-FP $x$.) If $x^{*}\neq x^{**}$, then $x^{*}$ is proper.
(ii) Consider the anchored right-regulated system, with frame $(-\infty,0]$. Consider the ML $x(\cdot)$, starting with the “empty” initial state $x(0)=x^{\emptyset}$. Then $x(t)$ is monotone non-increasing and $x(t)\to x^{*}$, where $x^{*}$ is an ML-FP, and moreover it is the maximum ML-FP. (That is, $x^{*}\geq x$ for any ML-FP $x$.) If $x^{*}\neq 1-x^{**}$, then $x^{*}$ is proper. (ML state $1-x^{**}$ corresponds to the entire particle mass at -$\infty$.)
@@
For an anchored two-side-regulated system, for any $v>0$, there exists a unique ML-FP, and the dependence on $v$ is continuous.
@@ 6 Preliminary definitions and results (generalizing those in [15]), related to SSAI
For a given $n$ consider the Markov process $x^{n}(\cdot)$ for an anchored system, and the corresponding centered process $\mathring{x}^{n}(\cdot)$. Process $\mathring{x}^{n}(\cdot)$ is not Markov, because the information about the location of the boundaries is “lost” after centering; it is a projection of Markov process $\mathring{x}^{n}(\cdot)$.
@@
Consider the anchored left-regulated system with frame $0,\infty)$. Suppose Assumptions 3.2, 3.3 and 3.4 hold. Then there exist $\bar{C}>0$ and $\bar{n}$ such that, uniformly in $n\geq\bar{n}$ and corresponding (for each $n$) speeds $v$ such that the process $x^{n}(\cdot)$ is stable, we have
@@
Suppose there exists a class $j$, $0\leq m_{j}<k_{j}$, and $\epsilon>0$, such that the component size distribution $F_{j}$ dominates a scaled (by a positive constant) measure such that $m_{j}$ components are zero and the remaining $d_{j}-m_{j}$ components distribution is the Lebesgue measure on $0,\epsilon]^{d_{j}-m_{j}}$. Then Assumption 6.2 holds.
@@
If the process $\mathring{x}^{n}(\cdot)$ in a free system is positive recurrent, then the average velocity $v_{n}$ of the unanchored free particle process $x^{n}(\cdot)$ is well-defined, with multiple possible equivalent definitions. For example,
$v_{n}=\lim_{t\to\infty}\Delta^{n}(t;\nu)/t,~~w.p.1,$ (19)
@@
Consider an anchored left-regulated system, with frame $0,\infty)$. We say that, in the anchored left-regulated system with frame $0,\infty)$, this condition holds for speed $v>0$, if for any sufficiently small $b\geq 0$, there is at most one DE-FP $x$ with $x_{\infty}=b$.
@@
Consider the left-regulated system with the frame $0,\infty)$. Suppose Assumptions 3.2, 3.3 and 3.4 hold. Consider our asymptotic regime, with the number of particles $n\to\infty$, such that for all large $n$ the process is stable and, moreover, $\rho^{n}=\mathbb{E}x^{n}_{0}(\infty)\to\rho<1$. Suppose that Condition 6.5 holds for speed $v>0$. Then SSAI holds for this $v$, i.e. $x^{n}(\infty)\Rightarrow x^{*}$, where $x^{*}$ is the ML-MFP and it is such that $x^{*}_{0}=\rho$.
@@
The proof of Lemma 6.8 is straightforward. Consider any finite locations of $m$-$\ell$ particles in the reduced system; consider original system with $m$-$\ell$ particles located as in the reduced system, $\ell$ particles located at C-C, $d_{j}-m$ particles located at $C$; let $C\uparrow\infty$.
@@ 7 Main results
Consider ML-FPs for a family of anchored systems, with the following set of parameters taking values within a compact set: speed $v$; limiting arrival rates $\{\sigma_{j}\}$; and the distributions $\{F_{j}\}$ of the job component sizes (convergence of distributions is in the weak sense). Assume further that the family of parameters is such that: speed $v$ is bounded away from 0; Assumption 3.2 holds uniformly; Assumption 3.3 holds uniformly in that it holds for the same class $j$, with the corresponding probability bounded away from 0, and $\sigma_{j}$ bounded away from 0. Then, for some $\bar{C}>0$, uniformly in that family of parameters and uniformly across all free and regulated systems, an ML-FP $x$ is such that $\Phi_{1}(\mathring{x})\leq\bar{C}<\infty.$
@@
Proof. For a two-side-regulated system, with frame $[0,B]$ there exists a unique ML-FP for any speed $v>0$. For a sequence of frame sizes $B\to\infty$, choose drift $v(B)$, such that the median of the fixed point is exactly in the middle, at point $B/2$. Let $x^{(B)}$ be such ML-FP shifted left by $B/2$, so that the median is at 0. Let us choose a subsequence of $B$ along which $v(B)\to v$. Limiting speed $v$ cannot be 0 or $\infty$, because otherwise the median could not stay in the middle. (Indeed, if the limiting speed is infinity, then the probability mass of the left boundary has to stay bounded away from 0 – otherwise the average drift to the right cannot match the average drift to the left; in fact, that mass goes to $1$, so the distribution concentrates on the left boundary. If the limiting speed is zero, then the distribution, in a weak limit, must concentrate at the right boundary – otherwise the average drift to the right cannot go to 0 to match the average drift to the left.) Then, by Theorem 7.1, $\Phi_{1}(\mathring{x}^{(B)})\leq C<\infty$, uniformly in all $c$. This in turn implies that $\Phi_{1}(x^{(B)})\leq C<\infty$, uniformly in all $B$, because $|\bar{x}^{(B)}|$ must stay bounded. So, we can choose a further subsequence of $B$ along which $x^{(B)}\to x$, where $x$ is proper. Then, by Lemma 5.12, $x$ is a proper free ML-FP with speed $v$. $\Box$
@@
$\alpha\int_{0}^{\infty}wdH(w)>v.$ (24)
Indeed, let $L(\beta)=\int_{0}^{\infty}e^{-\beta z}dH(z),~\beta\geq 0,$ be the Laplace transform of distribution $H(\cdot)$, and $\bar{L}(\beta)=\int_{0}^{\infty}e^{-\beta z}\bar{H}(z)dz=[1-L(\beta)]/\beta$, $\bar{H}(w)=1-H(w)$. Then $\chi^{(\epsilon)}$ is a DE-FP if
$\int_{-\infty}^{z}[-\chi^{(\epsilon)}_{w}]^{\prime}\alpha\bar{H}(z-w)dw=-[\chi^{(\epsilon)}_{z}]^{\prime}v,~~z\leq 0,$
@@
The LHS is convex, decreasing to 0 as $\beta\to\infty$, with $\bar{L}(0)=\int_{0}^{\infty}wdH(w)$, which is the mean component size. So, (25) uniquely determines $\beta$ as long as (24) holds.
Note that for any $0\leq\epsilon<1$,
@@
and $\chi^{(0)}$ is nothing else, but the stationary distribution of a particle moving in $(-\infty,0]$, with the rate $\alpha$ Poisson process of independent jumps to the right (but not over the right boundary at 0), with size distribution $H(\cdot)$, and moving left at speed $v$ between the jumps. We conclude that $\chi^{(\epsilon)}$ is the unique ML-FP (=ML-MFP) of the artificial $(\epsilon,0)$-reduced system with the right boundary at 0.
As will be seen in the proof of Theorem 7.3, $\chi^{(\epsilon)}$ gives the asymptotic shape of the left tail of any DE-FP with speed $v$. Informally, this is because, when $1-\epsilon$ is close to 0, the $(\epsilon,0)$-reduced original system is approximately the $(\epsilon,0)$-reduced artificial system.
@@
(i) In the anchored system, for any $v$ satisfying (24), there exists the unique DE-FP with the proper left tail. This solution will either (a) hit the $w$-axis or (b) will converge to some $\epsilon_{*}\geq 0$ as $w\uparrow\infty$. In case (a): for any $0\leq\epsilon<1$ there is a unique ML-FP with proper left tail for the $(\epsilon,0)$-reduced right-regulated system. In case (b): for any $\epsilon_{*}<\epsilon<1$ there is a unique ML-FP with proper left tail for the $(\epsilon,0)$-reduced right-regulated system; for $\epsilon=\epsilon_{*}$ there is a unique ML-FP with proper left and right tails for the $(\epsilon,0)$-reduced free system; for $\epsilon<\epsilon_{*}$ there is a unique ML-FP with proper left tail, but not proper on the right, for the $(\epsilon,0)$-reduced free system.
(ii) In the anchored system, for any $v$ satisfying (24), and any $0\leq\epsilon<1$, there exists at most one ML-FP with proper left tail for the $(\epsilon,0)$-reduced system with speed $v$; this ML-FP is either for the right-regulated system or for the free system, with proper or not proper right tail.
@@
(i) For the anchored free system, there exists a range $v_{min},v_{max}]$ of speeds $v$, with $0<v_{min}\leq v_{max}<\infty$ for which a proper ML-FP $x$ exists.